# Attractors for the Wave Equation with Nonlinear Energy Critical Sources and Dissipation

### v2

@@ Abstract
The wave equation with energy-critical nonlinear sources and dissipation in a bounded 3 D domain is considered. It is shown that the resulting dynamical system admits a global attractor. Under the additional assumption of strong monotonicity of the damping at the origin, the said attractor is also finite dimensional and smooth. In order to handle “energy criticality” of both sources and the damping, the methods used depend on enhanced dissipation [7], energy identity for weak solutions [21], an adaptation of Ball’s method [1], and the theory of quasi-stable systems [13].
@@
Keywords: Wave equation; global attractor; nonlinear damping; nonlinear source; critical exponents.
MSC: 35B33; 35B40; 35B41; 35L05; 35L70.
1 Introduction
@@
Let $\Omega\subset\mathbb{R}^{3}$, where $\Omega$ is a bounded domain with a sufficiently smooth boundary $\Gamma=\partial\Omega$. The following classical energy critical wave model is under consideration in this work:
@@
$\left\{\begin{aligned} &u_{tt}-\Delta u+g(u_{t})+f(u)=h,\quad\mbox{in}\quad\Omega\times\mathbb{R}^{+},\\ &u|_{\Gamma}=0,\quad\mbox{on}\quad\Gamma=\partial\Omega,\\ &u(x,0)=u_{0}(x),\quad u_{t}(x,0)=u_{1}(x),\quad x\in\Omega,\end{aligned}\right$. (1.1)
where $g$ is a nonlinear monotone damping of critical, quintic nature: $|g(s)|\leq M|s|^{5}$ and the source $f$ is also a nonlinear critical source with quintic upper bound: $|f(s)|\leq M|s|^{5},\,|s|>1$.
@@
A study of wave equation in 3 dimensions with a quintic source have attracted a lot of activities. Quintic term is critical for the wave equation in 3 dimensions, thus a host of mathematical issues at the level of well-posedness of weak solutions. In fact, while local well-posedness can be obtained via several methods, global solvability has been a source of major challenges. And here is the reason why. The strategy used for establishing existence of weak solutions to undamped wave equation relies heavily on Strichartz estimates which allow to trade differentiability into suitable integrability. However, Strichartz estimates are constructed for each solutions [so called Strichartz solution] $u$ on time interval $[0,T]$. These estimates loose uniformity with respect to bounded set of initial data when time $T\rightarrow\infty$, unless nonlinearity is subquintic. The above predicament, particularly sensitive to long time behavior of the models, stimulated enormous amount of research. A good account is presented in [20]. To address this issue, methods relaying on “nonconcentration” of singularity in S-estimates have been developed [9]. These allow to extend local solutions to global ones. Additional arguments relying on the so called backward regularity [29] lead to S-solutions which form a dynamical system. These methods also rely on finite speed of propagation which can be also proved in the presence of linearly bounded dissipation-[20]. However, in the case of highly nonlinear dissipation, proving globality and uniqueness of solutions by using the above described method appears challenging.
@@
This brings us to the topic of the present paper. We consider damping also of critical exponent with an enhanced dissipation. The problem goes back to [27] and [6, 7, 5] where such dynamics has been considered. In fact, the idea was to use the enhanced dissipation in order to establish well-posedness, particularly uniqueness of weak solutions corresponding to critical and supercritical sources.. The estimates in these references provide an initial road map to our analysis, which eventually culminates with the theory of coherent attracting sets. Thus, our goal is to take the problem to the next level by considering long time behavior-in particular existence and a structure of attracting sets. This will be accomplished by developing appropriate tools for establishing asymptotic smoothness [particularly in the degenerate case when $g^{\prime}(0)=0$], finite-dimensionality and smoothness of the attractors. The challenge here is due to double criticality of the source and dissipation. Criticality of the source makes potential energy functional $\int_{\Omega}F(u(x))dx$ with $F^{\prime}=f$ noncompact, thus preventing methods of compensated compactness from applicability. Criticality of the dissipation implies $g(u_{t})\in H^{-1}(\Omega)$, but $g(u_{t})\notin H^{-\alpha}(\Omega),\alpha<1$. The latter is the critical regularity for variational solutions to be well defined with test functions compactly embedded in $H^{1}(\Omega)$. Our final result establishes an existence of global attractor which, moreover, is finite-dimensional and smooth.
2 Functional Setting, Assumptions and Main Results
2.1 Functional Setting and the Assumptions Imposed
First of all, we will define some notations that will be used throughout the work. We denote by $H^{s}(\Omega)$ the $L^{2}$- based Sobolev’s space of the order $s$ and by $H^{s}_{0}(\Omega)$ the closure of $C^{\infty}_{0}(\Omega)$ in $H^{s}(\Omega)$. In particular, the inner products and norms in the spaces $L^{2}(\Omega)$ and $H^{1}_{0}(\Omega)$ will be denoted, respectively, by
$\left(u,v\right)=\int_{\Omega}uvdx,\quad\|u\|^{2}=\int_{\Omega}|u(x)|^{2}dx,$
$((u,v))=\left(\nabla u,\nabla v\right),\quad\|u\|^{2}_{H^{1}_{0}(\Omega)}=\|\nabla u\|^{2}.$
We consider the Laplacian operator $\Delta$ defined by the triplet $\{H^{1}_{0}(\Omega),L^{2}(\Omega),a(u,v)\}$ where $a(u,v)$ is the bilinear continuous form on $H^{1}_{0}(\Omega)\times H^{1}_{0}(\Omega)$ given by
$a(u,v)=\left(\nabla u,\nabla v\right),\quad u,v\in H^{1}_{0}(\Omega).$
It is known from Spectral Theorem for self-adjoint operators that the domain of $\Delta$ is given by $D[(-\Delta)]=H^{2}(\Omega)\cap H^{1}_{0}(\Omega)$ and there exists a complete orthonormal family of $L^{2}(\Omega)$, $\{\omega_{j}\}_{j\in\mathbb{N}}$, made of eigenfunctions of $\Delta$ under Dirichlet boundary conditions.
$\Delta\omega_{j}=\lambda_{j}\omega_{j},\;j=1,\cdots,\quad\mbox{and}\quad 0<\lambda_{1}\leq\lambda_{2}\leq\cdots,\lambda_{j}\rightarrow+\infty\quad\mbox{as}\quad j\rightarrow+\infty.$
We can also define the powers $(-\Delta)^{s}$ of $\Delta$, for $s\in\mathbb{R}$. The spaces
$H^{2s}=D[(-\Delta)^{s}],s\in\mathbb{R}\quad\text{ (with Dirichlet boundary conditions)}$
are Hilbert spaces with the norms $\|u\|_{H^{2s}}=\|(-\Delta)^{s}u\|$. We have $H^{2s_{1}}\subset H^{2s_{2}}$, $\forall s_{1},s_{2}\in\mathbb{R},$ $s_{1}\geq s_{2}$, where each space is dense in the following one and the injection is compact if $s_{1}>s_{2}$. Moreover
$\|(-\Delta)^{s_{2}}u\|\leq\lambda_{1}^{s_{2}-s_{1}}\|(-\Delta)^{s_{1}}u\|\quad\mbox{if}\quad s_{1}\geq s_{2}.$
We identify $D[(-\Delta)^{0}]=L^{2}(\Omega)$ with its dual, which leads to the identification $H^{-2s}=D[(-\Delta)^{-s}]$ with the dual space of $H^{2s}=D[(-\Delta)^{s}]$, $s>0$. In particular, we have $H^{0}=L^{2}(\Omega)$, $H^{1}=H^{1}_{0}(\Omega)$, and $H^{2}=H^{2}(\Omega)\cap H^{1}_{0}(\Omega)$. The operators $(-\Delta)^{s}$ are self-adjoints, i.e.,
$\left((-\Delta)^{s}u,v\right)=\left(u,(-\Delta)^{s}v\right),\quad\forall u,v\in H^{2s}.$
With the above notations we can define Hilbert spaces $\mathcal{H}_{s}:=H^{s+1}\times H^{s}$, endowed with the usual norms
[…]
In particular, with respect to the phase space corresponding to the weak solutions of problem (LABEL:P), the natural choice is
$\mathcal{H}:=\mathcal{H}_{0}=H^{1}\times H^{0}\equiv H^{1}_{0}(\Omega)\times L^{2}(\Omega),\\$
with norm given by
$U||^{2}_{\mathcal{H}}:=\|\nabla u\|^{2}+\|v\|^{2},\quad U\equiv(u,v).$
We assume the following hypotheses regarding the functions $h$, $g$ and $f$.
@@ 2.1 Functional Setting and the Assumptions Imposed
$h\in H^{0}$ is an external force and $f,g$ are given functions satisfying the following conditions:
@@
    $\kappa_{0}|s|^{4}\leq g^{\prime}(s)\leq\kappa_{1}\left(1+|s|^{4}\right),\quad\forall s\in\mathbb{R}.$ (2.1)
  • $f\in C^{2}(\mathbb{R})$ with $f(0)=0$ is of the following growth condition: there exists a positive constant $L_{f}>0$ such that
    $\left|f^{\prime\prime}(s)\right|\leq L_{f}\left(1+|s|^{3}\right),\quad\forall s\in\mathbb{R},$ (2.2)
    in conjunction with the dissipative condition
    $\lim_{|s|\rightarrow\infty}\frac{f(s)}{s}>-\lambda_{1}$ (2.3)
    where $\lambda_{1}$ is the first eigenvalue of the Laplacian operator $\Delta$ with Dirichlet boundary condition.
@@
$|f^{\prime}(s)|\leq C_{f}(1+|s|^{4}),\quad\forall s\in\mathbb{R},$ (2.4)
for some constant $C_{f}>0$. The dissipative condition (2.3) implies that there exists a constant $\nu\in[0,\lambda_{1})$ and $C_{\nu}>0$ such that
$-C_{\nu}-\frac{\nu}{2}|s|^{2}\leq F(s)\leq f(s)s+\frac{\nu}{2}|s|^{2},\quad\forall s\in\mathbb{R},$ (2.5)
where ${F(s):=\int_{0}^{s}f(\tau)d\tau}$.
@@
$\omega:=1-\frac{\nu}{\lambda_{1}}>0.$ (2.6)
@@
2.2 Statements of the Main Results
Of main interest to this paper are the so called “weak solutions”. There are several notions of “weak solutions” which are often refereed to as “finite energy solutions”. However, in the critical case one needs to be careful regarding precise meaning of weak solutions. While proving existence of strong [regular solutions] is rather routine, existence and the properties of weak solutions are more elusive in the critical case. Typical notion would be the so called “generalized” solutions which are defined as strong limits of semigroup regular solutions. Under certain conditions such solutions can be shown to satisfy variational form of equations-the so called “weak solutions”. Weak solutions are not necessarily limits of strong solutions [unless some compactness related properties are assumed]- so these are of interest to the study in “critical” situations. Here, by critical we refer to the property that potential energy is just $L^{1}$ integrable without any compactness properties assumed in prior literature [16, 27] and references therein. In what follows, we shall adopt rather classical definition -following [5, Definition 2.1] where solutions to (LABEL:P) under the Assumption 2.1 were treated.
@@ 2.2 Statements of the Main Results
$\begin{aligned} \int_{0}^{T}\int_{\Omega}\left[\,-u_{t}\psi_{t}+\nabla u\nabla\psi+g(u_{t})\psi+f(u)\psi\,\right]dxdt=-\left.\int_{\Omega}u_{t}\psi dx\right|_{0}^{T}+\int_{0}^{T}\int_{\Omega}h\psi dxdt,\end{aligned}$ (2.7)
@@
The first result refers to an existence and properties of weak solutions as defined in Definition 2.1.
@@
    $E(U(t))+\int_{s}^{t}\int_{\Omega}g(u_{t})u_{t}dxds=E(U(s)),\quad\mbox{for all}\quad t\geq s\geq 0,$ (2.8)
@@
    Hadamard well-posedness and stability: Let $U^{1}=(u^{1},u^{1}_{t})$ and $U^{2}=(u^{2},u^{2}_{t})$ be two weak solutions of problem (LABEL:P) with $U^{i}(0)=(u^{i}_{0},u^{i}_{1})=U^{i}_{0}\in\mathcal{H}$, $i=1,2$. Then
@@
The proof of the well-posedness theorem is based on the method introduced in [5]-which takes advantage of the enhanced dissipation exhibited by the additional integrability of the velocity. Existence of solutions was obtained in [7] by constructing suitable approximations of critical nonlinearities followed by the use of monotone operator theory. Uniqueness was shown in [5] by establishing first energy identity. However, continuous dependence stated in [5] holds for smoother initial data only. Since our ultimate goal is to have a properly defined dynamical system, we shall revisit the proof by using different method-based on Galerkin approximations. We will be able to establish full Hadamard well-posedness valid for all weak solutions. This is the main step toward construction of dynamical system $(\mathcal{H},S_{t})$, whose existence is guaranteed by Theorem 2.2. In addition, we shall show that all weak solutions can be obtained as strong limits of Galerkin solutions. This is in line with [20] and allows to use Galerkin approximation in the estimates where there is not enough of smoothness to use operational calculus and classical limiting arguments.
@@
The main results of this work describe long-time behavior of solutions generated by the dynamical system $(\mathcal{H},S_{t})$. The definitions of the respectful quantities along with the known properties are relegated, for the convenience of a reader, to the Appendix 7.
@@
$\lim_{t\rightarrow-\infty}\mbox{dist}_{\mathcal{H}}(U(t),\mathcal{N})=0\quad\mbox{and}\quad\lim_{t\rightarrow+\infty}\mbox{dist}_{\mathcal{H}}(U(t),\mathcal{N})=0.$
The proof of the existence of a global attractor is based on an adaptation of “Ball’s” method that employs energy arguments and is fundamentally based on the validity of energy equality valid for weak solutions. This method is adapted to the present context by incorporating a nonlinear dissipation. The latter is possible due to monotonicity properties.
Under the additional assumptions imposed on the damping $g^{\prime}(0)>0$, the following properties of the said attractor $\mathfrak{A}$ are valid.
@@
[Quasi-stability property] Let the Assumption (2.1) holds and $g^{\prime}(0)>0$. Then the dynamical system $(\mathcal{H},S_{t})$ is quasi-stable.
The proof that the dynamical system is quasi-stable follows from the fact that the dynamical system $(\mathcal{H},S_{t})$ satisfies the so called quasi-stability inequality, which is an enhanced version of squeezing property -see the Appendix 7.
@@
    Moreover, there exists $R_{*}>0$ such that
    $\|u_{tt}(t)\|^{2}+\|u_{t}(t)\|^{2}_{H^{1}}+\|u(t)\|^{2}_{H^{2}}\leq R^{2}_{*},$
    for all $t\in\mathbb{R}$ and some positive constant $R_{*}$.
@@
The result presented above are inspired by earlier papers [5, 7] where the model with critical [quintic] sources and enhanced dissipations were considered. For these models an existence and uniqueness of weak solutions has been established in references cited above. Hadamard well-posedness has been shown in [5] for subcritical sources only. Questions related to blow up of weak solutions have been investigated in [4, 6, 8, 15]. Full Hadamard well-posedness and complete analysis of long time dynamics is the main contribution of the present manuscript.
@@
We conclude this section by briefly outlining general approach taken in the paper. The existence of weak solutions is established by Galerkin method with suitable $L^{p}$ global estimates. Weak solutions are shown to satisfy energy equality. Here, again, enhanced dissipation plays a major role. This is a critical step in order to prove uniqueness of weak solutions and Hadamard well-posedness. The latter allows to construct a well-posed dynamical system. By resorting, again, to enhanced dissipation, the system $(\mathcal{H},S_{t})$ is shown to be ultimately dissipative, hence admitting weak global attractor. The next critical step is asymptotic compactness. Relying on already proved energy identity, an adaptation of the method of [1] to account for nonlinear damping, allows to establish asymptotic compactness, hence an existence of global attractor. Under the additional assumption of non-degeneracy of the damping at the origin, global attractor is shown to be: smooth and also finite-dimensional. In addition, the dynamical system has also exponential attractor. These last properties follow from the established quasi-stability inequality.
2.3 Past Contributions.
@@
Wave dynamics in 3 dimensions with subquintic and quintic source has been an object of long-time and intense activities. The main issues of interest to this work are well-posedness and long time behavior. The presence of the quintic source makes the traditional methodology un-applicable-due to the lack of appropriate embeddings which would “close” the needed estimates. Relatively recent development of Strichartz estimates opened the door to novel methodologies, where the additional space integrability is achieved due to compensation with time integrability. Careful balance of time-space integrability allows to construct local solutions referred to as Shatah-Struwe [S-S] solutions. However, such solutions do not admit global extensions [unless the case is subcritical]. Here, the rescue comes by appealing to non-concentration argument [9]. This argument relies on finite speed of propagation, hence sensitive to nonlinear damping. And this is one of the main points where nonlinearity of the damping in the present model becomes predicament. In fact, in the case of linearly bounded dissipation, the methodology pursued in [20]. In the case of critical sources and linear damping is the most effective. The starting point of the analysis in [20] is local existence of S-S-solutions. These are obtained by using Strichartz estimates applied to Galerkin solutions. Since Strichartz estimates do not admit global extensions [unless the case is subcritical], global existence of Galerkin limits is obtained by appealing to non-concentration argument [9]. Having established global existence [S-estimates do not blow up in time]-the next step is a construction of a dynamical system which contains weak Galerkin limits [not necessarily S-solutions which may be lost in the limiting process]. The latter requires uniqueness and continuous dependence of the said solutions. In order to achieve this, the method of backward trajectories is employed in order to generate smoothness of backward trajectories. Classical forward regularity argument allows to propagate the smoothness to any time $T$. This allows to claim uniqueness of orbits and also energy equality on any time interval. Equipped with the above properties, energy method is used in [20] in order to assert an asymptotic compactness and to conclude the proof of existence of global attractor. The orbits in the attractor become eventually S-S solutions.

### v4

# Stabilizing energy-critical wave equation to a finite dimensional attractor via nonlinear damping

@@ Abstract.
The wave equation with energy critical sources and nonlinear damping defined on a 3D bounded domain is considered. It is shown that the resulting dynamical system admits a global attractor. Under the additional assumption of strong monotonicity of the damping at the origin, it is shown that the originally unstable quintic wave is uniformly stabilised to a finite dimensional and smooth set. Moreover, the existence of exponential attractor is established. In order to handle “energy criticality” of both sources and damping, the methods used depend on enhanced dissipation [7], energy identity for weak solutions [23], an adaptation of Ball’s method [1], and the theory of quasi-stable systems [15].
Mathematical Subject Classification 2020: 35B33; 35B40; 35B41; 35L05; 35L70.
@@
Keywords: Wave equation with quintic source, finite dimensional attractor; nonlinear damping.
Research partially supported by National Science Foundation: NSF-DMS- 2205508.
Research partially supported by Fundect/CNPq Grant 15/2024.
1. Introduction
@@
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded domain with a sufficiently smooth boundary $\Gamma=\partial\Omega$. The following classical energy critical wave model is under consideration in this work:
@@
(1.1) $\left\{\begin{aligned} &u_{tt}-\Delta u+g(u_{t})+f(u)=h,\quad\mbox{in}\quad\Omega\times\mathbb{R}^{+},\\ &u|_{\Gamma}=0,\quad\mbox{on}\quad\Gamma=\partial\Omega,\\ &u(x,0)=u_{0}(x),\quad u_{t}(x,0)=u_{1}(x),\quad x\in\Omega,\end{aligned}\right$.
where $g$ is a nonlinear monotone damping of critical, quintic nature: $|g(s)|\leq M|s|^{5}$ and the source $f$ is also a nonlinear critical source with quintic upper bound: $|f(s)|\leq M|s|^{5},\,|s|>1$. Quintic wave equations, which include nonlinear terms to the fifth power, are used to model complex physical phenomena like solitons in optics and plasma physics, vibrations in materials, fluid dynamics, and the dynamics of wave propagation in various media. They are crucial for understanding and predicting the behavior of nonlinear systems and help in designing technologies in fiber optics, plasma confinement, solid state physics and material science [molecular vibrations and wave propagation in harmonic crystals],and nonlinear circuits.
@@
A study of wave equation in 3 dimensions with a quintic source have attracted a lot of activities. Quintic term is energy-critical for the wave equation in 3 dimensions, thus a host of mathematical issues starting with well-posedness of weak solutions. While local well-posedness can be obtained via several methods, global solvability has been a source of major challenges. And here is the reason why. The strategy used for establishing existence of weak solutions to undamped wave equation relies heavily on Strichartz estimates which allow to trade differentiability into suitable integrability. However, Strichartz estimates are constructed for each solution [so called Strichartz solution] $u$ on time interval $[0,T]$. These estimates loose uniformity with respect to bounded set of initial data when time $T\rightarrow\infty$, unless nonlinearity is subquintic. The above predicament, particularly sensitive to long time behavior of the models, stimulated great amount of research -see [22]. To address this issue, methods relaying on “nonconcentration” of singularity in S-estimates have been developed [10]. These allow to extend local solutions to global ones. Additional arguments relying on the so called backward regularity [32] lead to S-solutions forming a dynamical system which eventually is shown to have a global attractor. These methods rely on finite speed of propagation which can be also proved in the presence of linearly bounded dissipation-[22]. However, in the case of highly nonlinear dissipation, proving globality and uniqueness of solutions by using the above described method appears challenging.
@@
This brings us to the topic of the present paper. Our aim is, in addition to the study of attracting sets, to construct stabilizing effect of nonlinear damping which may eventually bring [in the uniform topology of the phase space] the originally unstable hyperbolic dynamics to a finite dimensional configuration -characterised by a compact set-attractor. Such property is of critical value within the realm of control theory, where finite dimensional control methods could be applied to forge a desirable outcome for the system. To accomplish this, we consider damping also of critical exponent with an enhanced dissipation effect. The dea goes back to [30] and [6, 7, 5] where such dynamics has been considered. In fact, the enhanced dissipation was used in order to establish well-posedness, particularly uniqueness of weak solutions corresponding to critical and supercritical sources.. The estimates in these references provide an initial road map to our analysis, which eventually culminates with the theory of coherent attracting sets. Thus, our goal is to take the problem to the next level by considering long time behavior-in particular existence and a structure of attracting sets including finite dimensionality and an existence of exponentially attracting sets. This will be accomplished by developing appropriate tools for asymptotic smoothness [particularly in the degenerate case when $g^{\prime}(0)=0$], finite-dimensionality and maximal regularity of the attractors. The challenge here is due to double criticality of the source and the dissipation. Criticality of the source makes potential energy functional $\int_{\Omega}F(u(x))dx$ with $F^{\prime}=f$ noncompact, thus preventing methods of compensated compactness from applicability. Criticality of the dissipation implies $g(u_{t})\in H^{-1}(\Omega)$, but $g(u_{t})\notin H^{-\alpha}(\Omega),\alpha<1$. The latter is the critical regularity for variational solutions to be well defined with compact test functions in $H^{1}(\Omega)$. Our final result provides an existence of global attractor which, moreover, is finite-dimensional and smooth along with an existence of exponential attractor is also established. Thus, quintic waves are stabilised uniformly/exponentially to finite dimensional compact sets.
2. Functional Setting, Assumptions and Main Results
2.1. Functional Setting and the Assumptions Imposed
First of all, we will define some notations that will be used throughout the work. We denote by $H^{s}(\Omega)$ the $L^{2}$- based Sobolev’s space of the order $s$ and by $H^{s}_{0}(\Omega)$ the closure of $C^{\infty}_{0}(\Omega)$ in $H^{s}(\Omega)$. We also denote $\left(u,v\right)=(u,v)_{L^{2}(\Omega)}$ and $\|u\|^{2}=\|u\|^{2}_{L^{2}(\Omega)}$ for the inner product and norm in $H^{0}:=L^{2}(\Omega)$, and $((u,v))=\left(\nabla u,\nabla v\right)$ and $\|u\|^{2}_{H^{1}_{0}(\Omega)}=\|\nabla u\|^{2}$ for those in $H^{1}:=H^{1}_{0}(\Omega)$. We consider the Laplacian operator $\Delta$ defined by the triplet $\{H^{1},H^{0},a(u,v)\}$ where $a(u,v)$ is the bilinear continuous form on $H^{1}\times H^{1}$ given by $a(u,v)=\left(\nabla u,\nabla v\right)$, $u,v\in H^{1}.$ With respect to the phase space corresponding to the weak solutions of problem (LABEL:P), our analysis is carried out in the Sobolev space $\mathcal{H}:=\mathcal{H}_{0}=H^{1}\times H^{0},$ equipped with the standard norm $U||^{2}_{\mathcal{H}}:=\|\nabla u\|^{2}+\|v\|^{2},\quad U\equiv(u,v).$
We assume the following hypotheses imposed on the functions $h$, $g$ and $f$.
@@ 2.1. Functional Setting and the Assumptions Imposed
$h\in H^{0}$ - an external force, $f,g$ are subject to the following conditions:
@@
    (2.1) $\kappa_{0}|s|^{4}\leq g^{\prime}(s)\leq\kappa_{1}\left(1+|s|^{4}\right),\quad\forall s\in\mathbb{R}.$
  • $f\in C^{2}(\mathbb{R})$, $f(0)=0$ and there exists a positive constant $L_{f}>0$ such that
    (2.2) $\left|f^{\prime\prime}(s)\right|\leq L_{f}\left(1+|s|^{3}\right),\quad\forall s\in\mathbb{R}.$
    The following standard dissipative condition is imposed.
    (2.3) $\lim_{|s|\rightarrow\infty}\frac{f(s)}{s}>-\lambda_{1}$
    where $\lambda_{1}$ is the first eigenvalue of the Laplace operator $\Delta$ with Dirichlet boundary data.
@@
(2.4) $|f^{\prime}(s)|\leq C_{f}(1+|s|^{4}),\quad\forall s\in\mathbb{R},$
for some constant $C_{f}>0$. Condition (2.3) implies that there exists a constant $\nu\in[0,\lambda_{1})$ and $C_{\nu}>0$ such that $F(s):=\int_{0}^{s}f(\tau)d\tau$ satisfies:
(2.5) $-C_{\nu}-\frac{\nu}{2}|s|^{2}\leq F(s)\leq f(s)s+\frac{\nu}{2}|s|^{2},\quad\forall s\in\mathbb{R},$
@@
(2.6) $\omega:=1-\frac{\nu}{\lambda_{1}}>0.$
@@
2.2. Statements of the Main Results
Of main interest to this paper are the so called “weak solutions”. There are several notions of “weak solutions” which are often refereed to as “finite energy solutions”. However, in the critical case one needs to be careful regarding precise meaning of weak solutions. While proving existence of strong [regular solutions] is rather routine, existence and the properties of weak solutions are more elusive in the critical case. Typical notion would be the so called “generalized” solutions which are defined as strong limits of semigroup regular solutions. Under certain conditions, such solutions can be shown to satisfy variational form of equations-the so called “weak solutions”. Weak solutions are not necessarily limits of strong solutions [unless some compactness related properties are assumed]- so these are of interest to the study in “critical” situations. Here, by critical we refer to the property that potential energy is just $L^{1}$ integrable without any compactness properties assumed in prior literature [18, 30] and references therein. In what follows, we shall adopt rather classical definition -following [5, Definition 2.1] where solutions to (LABEL:P) under the Assumption 2.1 were treated.
@@ 2.2. Statements of the Main Results
(2.7) $\int_{0}^{T}\int_{\Omega}\left[\,-u_{t}\psi_{t}+\nabla u\nabla\psi+g(u_{t})\psi+f(u)\psi\,\right]dxdt=-\left.\int_{\Omega}u_{t}\psi dx\right|_{0}^{T}+\int_{0}^{T}\int_{\Omega}h\psi dxdt,$
@@
The first result refers to an existence and properties of weak solutions-defined in Definition 2.1.
@@
    (2.8) $E(U(t))+\int_{s}^{t}\int_{\Omega}g(u_{t})u_{t}dxds=E(U(s)),\quad\mbox{for all}\quad t\geq s\geq 0,$
@@
    Hadamard well-posedness : Let $U^{1}=(u^{1},u^{1}_{t})$ and $U^{2}=(u^{2},u^{2}_{t})$ be two weak solutions of problem (LABEL:P) with $U^{i}(0)=(u^{i}_{0},u^{i}_{1})=U^{i}_{0}\in\mathcal{H}$, $i=1,2$. Then
@@
The proof of the well-posedness theorem is based on the method introduced in [5]-which takes advantage of the enhanced dissipation exhibited by the additional integrability of the velocity. Existence of solutions was obtained in [7] by constructing suitable approximations of critical nonlinearities followed by the use of monotone operator theory. Uniqueness was shown in [5] by establishing first energy identity. However, continuous dependence stated in [5] holds for smoother initial data only. Since our ultimate goal is to have a properly defined dynamical system, we shall revisit the proof by using different method-based on Galerkin approximations. We will be able to establish full Hadamard well-posedness valid for all weak solutions in the energy critical case. This is the main step toward construction of dynamical system $(\mathcal{H},S_{t})$, whose existence is guaranteed by Theorem 2.2. In addition, we shall show that all weak solutions can be obtained as strong limits of Galerkin solutions. This is in line with [22] and allows to use Galerkin approximation in the estimates where there is not enough of smoothness to use operational calculus and classical limiting arguments.
@@
The main results of this work describe long-time behavior of solutions generated by the dynamical system $(\mathcal{H},S_{t})$. To maintain the text concise and avoid overloading it with preliminaries, the definitions and results related to the abstract theory of dynamical systems—particularly those concerning global attractors and their properties—will be referenced to the results established in established sources. For the sake of concreteness, we shall provide precise references to [15, 16].
@@
(2.11) $\lim_{t\rightarrow-\infty}\mbox{dist}_{\mathcal{H}}(U(t),\mathcal{N})=0\quad\mbox{and}\quad\lim_{t\rightarrow+\infty}\mbox{dist}_{\mathcal{H}}(U(t),\mathcal{N})=0.$
The proof of existence of global attractor is based on generalization of “Ball’s” method that employs energy arguments and is fundamentally based on the validity of energy equality valid for weak solutions. This method is adapted to the present context by establishing first the validity of energy equality and by incorporating critical nonlinear dissipation. The latter is possible due to the monotonicity properties.
Under the additional assumption imposed on the damping $g^{\prime}(0)>0$, the following properties of the attractor $\mathfrak{A}$ are valid.
@@
[Quasi-stability property.] Let the Assumption (2.1) holds and $g^{\prime}(0)>0$. Then, the dynamical system $(\mathcal{H},S_{t})$ is quasi-stable.
The proof that the dynamical system is quasi-stable follows from the fact that the dynamical system $(\mathcal{H},S_{t})$ satisfies the so called quasi-stability inequality, which is an enhanced version of squeezing property -see [16, Definition 7.9.2].
@@
    Moreover, there exists $R_{*}>0$ such that: $\|u_{tt}(t)\|^{2}+\|u_{t}(t)\|^{2}_{H^{1}}+\|u(t)\|^{2}_{H^{2}}\leq R^{2}_{*},$ for all $t\in\mathbb{R}$ and some positive constant $R_{*}$.
@@
The result presented above are inspired by earlier papers [5, 7] where the model with critical [quintic] sources and enhanced dissipations were considered. For these models an existence and uniqueness of weak solutions has been established in references cited above. Hadamard well-posedness has been shown in [5] for subcritical sources only. Questions related to blow up of weak solutions have been investigated in [4, 6, 9, 17]. Full Hadamard well-posedness and complete analysis of long time dynamics [including finite-dimensionality of the attractor] is the main contribution of the present manuscript.This includes the final assertion that quintic waves can be stabilized [uniformly/exponetially] to a finite-dimensional set and characterised by finitely many degrees of freedom. This opens the door to a construction of determining functionals [15, 24] and data assimilation methods which provide a basis for computational analysis.
@@
We conclude this section by briefly outlining general approach taken in the paper. The existence of weak solutions is established by Galerkin method with suitable $L^{p}$ global estimates. Weak solutions are shown to satisfy energy equality. Here, again, enhanced dissipation plays a major role. This is a critical step in order to prove uniqueness of weak solutions and Hadamard well-posedness. The latter allows to construct a well-posed dynamical system. By resorting, again, to enhanced dissipation, the system $(\mathcal{H},S_{t})$ is shown to be ultimately dissipative, hence admitting weak global attractor. The next critical step is asymptotic compactness. Relying on already proved energy identity, an adaptation of the method of [1] to account for nonlinear damping, allows to establish asymptotic compactness, hence an existence of global attractor. Under the additional assumption of non-degeneracy of the damping at the origin, global attractor is shown to be: smooth and also finite-dimensional. In addition, the dynamical system has also exponential attractor. These last properties follow from the proven quasi-stability inequality.
2.3. Past contributions-history of the problem.
@@
Wave dynamics in 3 dimensions with subquintic and quintic source has been an object of long-time and intense activities. The main issues of interest to this work are well-posedness and long time behavior. The presence of the quintic source makes the traditional methodology un-applicable-due to the lack of appropriate embeddings which would “close” the needed estimates. Relatively recent development of Strichartz estimates opened the door to novel methodologies, where the additional space integrability is achieved due to compensation with time integrability. Careful balance of time-space integrability allows to construct local solutions referred to as Shatah-Struwe [S-S] solutions. However, such solutions do not admit global extensions [unless the case is subcritical]. Here, the rescue comes by appealing to non-concentration argument [10]. This argument relies on finite speed of propagation, hence sensitive to nonlinear damping. And this is one of the main points where nonlinearity of the damping in the present model becomes predicament. In fact, in the case of linearly bounded dissipation, the methodology pursued in [22] is effective. The starting point of the analysis in [22] is local existence of S-S-solutions. These are obtained by using Strichartz estimates applied to Galerkin solutions. Since Strichartz estimates do not admit global extensions [unless the case is subcritical], global existence of Galerkin limits is obtained by appealing to non-concentration argument [10]. Having established global existence [S-estimates do not blow up in time]-the next step is a construction of a dynamical system which contains weak Galerkin limits [not necessarily S-solutions which may be lost in the limiting process]. The latter requires uniqueness and continuous dependence of the said solutions. In order to achieve this, the method of backward trajectories is employed in order to generate smoothness of backward trajectories. Classical forward regularity argument allows to propagate the smoothness to any time $T$. This allows to claim uniqueness of orbits and also [due to the established backward smoothness] energy equality on any time interval. Equipped with the above properties, energy method is used in [22] in order to assert an asymptotic compactness and to conclude the proof of existence of global attractor. The orbits in the attractor become eventually S-S solutions.